# Optimistic Multi-Agent Policy Gradient

### v2

@@ Abstract
Relative overgeneralization (RO) occurs in cooperative multi-agent learning tasks when agents converge towards a suboptimal joint policy due to overfitting to suboptimal behavior of other agents. No methods have been proposed for addressing RO in multi-agent policy gradient (MAPG) methods although these methods produce state-of-the-art results. To address this gap, we propose a general, yet simple, framework to enable optimistic updates in MAPG methods that alleviate the RO problem. Our approach involves clipping the advantage to eliminate negative values, thereby facilitating optimistic updates in MAPG. The optimism prevents individual agents from quickly converging to a local optimum. Additionally, we provide a formal analysis to show that the proposed method retains optimality at a fixed point. In extensive evaluations on a diverse set of tasks including the Multi-agent MuJoCo and Overcooked benchmarks, our method outperforms strong baselines on 13 out of 19 tested tasks and matches the performance on the rest.
@@ 1 Introduction
Multi-agent reinforcement learning (MARL) is a promising approach for many cooperative multi-agent decision making applications, such as those found in robotics and wireless networking (Busoniu et al., 2008). However, despite recent success on increasingly complex tasks (Vinyals et al., 2019; Yu et al., 2022), these methods can still fail on simple two-player matrix games as shown in Figure 1. The underlying problem is that in cooperative tasks, agents may converge to a suboptimal joint policy when updating their individual policies based on data generated by other agents’ policies which have not converged yet. This phenomenon is called relative overgeneralization (Wiegand, 2004) (RO) and has been widely studied in tabular matrix games (Claus & Boutilier, 1998; Lauer & Riedmiller, 2000; Panait et al., 2006) but remains an open problem in state-of-the-art MARL methods (De Witt et al., 2020; Yu et al., 2022; Kuba et al., 2021).
@@
The cause for RO can be understood intuitively. In a cooperative multi-agent system, the common reward derives from the joint actions. From the perspective of a single agent, an optimal individual action may still incur low joint reward due to the non-cooperative behavior of other agents. This is common in cooperative tasks when iteratively optimizing individual policies, especially at the beginning of training since individual agents have not learned to act properly to cooperate with others. It is particularly challenging in tasks with a large penalty for incorrect joint actions, such as the pitfalls in the climbing matrix game in Figure 1 leading often to agents that prefer a suboptimal joint policy.
@@
Existing techniques to overcome the pathology share the idea of applying optimistic updating in Q-learning with different strategies to control the degree of optimism (Lauer & Riedmiller, 2000) such as in hysteretic Q-learning (Matignon et al., 2007) or in lenient agents (Panait et al., 2006). However, these methods are designed based on Q-learning and only tested on matrix games (Matignon et al., 2012) with tabular Q representation. Although optimism successfully helps overcome RO problem and converges to global optima in tabular tasks, it could amplify the overestimation problem when combined with DQN (Van Hasselt et al., 2016) with function approximation (Omidshafiei et al., 2017; Palmer et al., 2017; Rashid et al., 2020a), as verified in our experiments. On the other hand, recent multi-agent policy gradient (MAPG) methods have achieved state-of-the-art performance on popular MARL benchmarks (De Witt et al., 2020; Yu et al., 2022; Sun et al., 2022; Wang et al., 2023) but still suffer from RO problem, converging to a low value local optimum in simple matrix games. The above drawbacks motivate this work to investigate whether optimism can be applied to MAPG methods and whether it can boost performance further by mitigating the RO problem.
@@
Our contribution is threefold: (1) To our knowledge, we are the first to investigate the application of optimism in MAPG methods. We propose a general, yet simple, framework to incorporate optimism into the policy gradient computation in MAPG. Specifically, we propose to clip the advantage values when updating the policy. For completeness, we extend the framework to include a hyperparameter to control the degree of optimism, resulting in a Leaky ReLU function (Maas et al., 2013) to reshape the advantage values. (2) We provide a form analysis to show that the proposed method retains optimality at a fixed point. (3) In our experiments111Source Code: https://github.com/wenshuaizhao/optimappo, the proposed OptiMAPPO algorithm successfully learns global optima in matrix games and outperforms both recent state-of-the-art MAPG methods MAPPO (Yu et al., 2022), HAPPO, HATRPO (Kuba et al., 2021), and existing optimistic methods in complex domains.
@@ 2 Related work
In this section, we discuss classic optimistic methods and optimistic DQN based approaches. We also discuss recent MAPG methods that yield state-of-the-art performance on common benchmarks. Optimistic Thompson sampling (O-TS) (Hu et al., 2023) is also discussed as it utilizes a similar clipping technique to improve the exploration for stochastic bandits. General multi-agent exploration methods are related and introduced. For completeness, we further discuss advantage shaping in single-agent settings.
@@
To the best of our knowledge, distributed Q-learning (Lauer & Riedmiller, 2000) proposes the first optimistic updating in independent Q-learning, where -values are only updated when the value increases. (Lauer & Riedmiller, 2000) also provides brief proof that the proposed optimism based method converges to the global optimum in deterministic environments. To handle stochastic settings, hysteretic Q-learning (Matignon et al., 2007) adjusts the degree of optimism by setting different learning rates for the Q values of actions with different rewards. The frequency maximum Q value heuristic (FMQ) (Kapetanakis & Kudenko, 2002) considers changing the action selection strategies during exploration instead of modifying the updating of Q values. Lenient agent (Panait et al., 2006; Wei & Luke, 2016) employs more heuristics to finetune the degree of optimism by initially adopting an optimistic disposition and gradually transforming into average reward learners.
@@
Recently, several works have extended optimistic methods and lenient agents to Deep Q-learning (DQN). Dec-HDRQN (Omidshafiei et al., 2017) and lenient-DQN (Palmer et al., 2017) apply hysteretic Q-learning and leniency to DQN, respectively. Optimistic methods have also been combined with value decomposition methods. Weighted QMIX (Rashid et al., 2020a) uses a higher weight to update the Q-values of joint actions with high rewards and a lower weight to update the values of suboptimal actions. FACMAC (Peng et al., 2021) improves MADDPG (Lowe et al., 2017) by taking actions from other agents’ newest policies while computing the Q values, which can also be regarded as optimistic updating as it is natural to suppose the newer policies of other agents would generate better joint Q-values. Even though the optimism has been applied to DQN-based methods, in common benchmarks they are usually outperformed by recent MAPG methods (Yu et al., 2022). This unsatisfying performance of optimistic DQN methods can be attributed to the side effect of overestimation of Q-values, which we also empirically verify in our experiments.
@@
COMA (Foerster et al., 2018) is an early MAPG method in parallel with value decomposition methods (Sunehag et al., 2017; Rashid et al., 2020b). It learns a centralized on-policy function and utilizes it to compute the individual advantage for each agent to update the policy. However, until IPPO and MAPPO (De Witt et al., 2020; Yu et al., 2022) which directly apply single-agent PPO (Schulman et al., 2017) into multi-agent learning, MAPG methods show significant success on popular benchmarks. The strong performance of these methods might be credited to the property that individual trust region constraint in IPPO/MAPPO can still lead to a centralized trust region as in single-agent PPO (Sun et al., 2022). Nonetheless, HAPPO/HATRPO (Kuba et al., 2021) and A2PO (Wang et al., 2023) further improve the monotonic improvement bound by enforcing joint and individual trust region, while with the cost of sequentially updating each agent. However, even with a strong performance on popular benchmarks, these methods don’t explicitly consider the relative overgeneralization problem during multi-agent learning and can still converge to a suboptimal joint policy.
@@
Since RO can be seen as a special case of the general exploration problem we discuss below general MARL exploration methods. Similar to single-agent exploration, (Rashid et al., 2020b; Hu et al., 2021) use noise for exploration in multi-agent learning. However, as shown by (Mahajan et al., 2019) noise-based exploration can result in suboptimal policies. For coordinated exploration, multi-agent variational exploration (Mahajan et al., 2019) conditions the joint Q-value on a latent state. (Jaques et al., 2019) maximizes the mutual information between agent behaviors for coordination. This may nevertheless still lead to a sub-optimal joint strategy (Li et al., 2022). (Zhao et al., 2023) propose conditionally optimistic exploration (COE) which augments agents’ Q-values by an optimistic bonus based on a global state-action visitation count of preceding agents. However, COE is designed for discrete states and actions and is difficult to scale up to complex tasks with continuous state and action spaces. Cooperative multi-agent exploration (CMAE) (Liu et al., 2021) only counts the visitations of states in a restricted state space to learn an additional exploration policy. However, the restricted space selection is hard to scale up and thus CMAE fails to show performance improvement on widely used MARL benchmarks. Our method aims to mitigate the specific RO problem with state-of-the-art MAPG methods in order to further boost their performance on complex tasks.
@@ 3.1 Problem Formulation
We mainly study the fully cooperative multi-agent sequential decision-making tasks which can be formulated as a decentralized Markov decision process (Dec-MDP) (Bernstein et al., 2002) consisting of a tuple , where is the set of agents. At time step of Dec-MDP, each agent observes the full state in the state space of the environment, performs an action in the action space from its policy . The joint policy consists of all the individual policies . The environment takes the joint action of all agents , changes its state following the dynamics function and generates a common reward for all the agents. is a reward discount factor. The agents learn their individual policies and maximize the expected return: .
@@ 4.1 Optimistic MAPPO
The advantage is usually estimated by the generalized advantage estimator (GAE) defined as
@@
In order to overcome the relative overgeneralization problem in MAPG methods, the proposed optimistic MAPPO (OptiMAPPO) applies a clipping operation to reshape the estimated advantages (Schulman et al., 2015). OptiMAPPO optimizes the agents’ policies by maximizing the following new objective
@@
The proposed clipping operation can be seen as a special case of a Leaky ReLU (LR) operation of advantage values where there is a hyperparameter to control the degree of optimism, . Our clipping operation is the case when , while recovers the original MAPG methods. In our experiments, we find that the performance is improved more while setting lower , i.e. higher optimism. However, we argue that such an extension could be beneficial in stochastic reward environments as discussed in hysteretic Q-learning (Matignon et al., 2007) and lenient agents methods (Palmer et al., 2017; Matignon et al., 2012). The extension allows us to control the degree of optimism in a finer granularity. We leave the study in stochastic environments as future work and this paper is primarily focused on the first to investigate the effectiveness of optimism in MAPG methods.
@@ 5.1 Repeated Matrix Games
Even though with small state and action spaces, the climbing and penalty matrix games (Claus & Boutilier, 1998), as shown in Figure 1, are usually hard to obtain the optimal joint solution without explicitly overcoming the relative overgeneralization problem. The matrix games have two agents which select the column and row index of the matrix respectively. The goal is to select the correct row and column index to obtain the maximal element of a matrix. In the penalty game, is the penalty term and we evaluate for . The lower value of , the harder for agents to identify the optimal policy due to the growing risk of penalty. Following (Papoudakis et al., 2020), we set constant observation and the episode length of repeated games as .
@@ 5.6 Does Q-learning Based Optimism Work?
We empirically analyze whether Q-learning based optimism can boost performance in the Overcooked tasks using hysteretic Q-learning (Omidshafiei et al., 2017; Palmer et al., 2017). Figure 5 shows average Q values during training with different degrees of optimism. When decreases excessively, i.e. optimism increases too much, performance decreases. In both tasks shown in Figure 5, the highest optimism estimates the highest Q-values while showing the worst performance.
@@ 6 Limitation and Conclusion
Our optimistic updating approach yields state-of-the-art performance. However, as with other optimistic updating methods (Lauer & Riedmiller, 2000; Matignon et al., 2007), optimism can lead to a sub-optima when misleading stochastic rewards exist. The proposed Leaky ReLU extension allows further adaptive adjustment of the optimism degree to balance between optimism and neutrality and may allow to reduce the severity of stochastic rewards but this requires future investigation. In addition, lenient agents (Panait et al., 2006; Wei & Luke, 2016; Palmer et al., 2017) provide a set of heuristic techniques to adapt the degree of optimism, applicable to our method, to mitigate the problem with stochastic rewards. We leave this also as future work.
@@ C.2 Key Hyper-Parameters
In all the tasks of MA-MuJoCo, we use the same hyperparameters listed in Table 5. The implementation is based on the HAPPO (Kuba et al., 2021) codebase, and the other hyperparameters are the default.
@@ Appendix D More Results of Popular Deep MARL methods on Matrix Games
In order to show the RO problem in existing MARL methods, we cite the following results of common deep MARL methods on the penalty and climbing games from the benchmarking paper (Papoudakis et al., 2020). As shown in Table 7, popular MARL methods without explicitly considering the RO problem fail to solve the matrix games.
@@ Appendix E Results of Optimistic MAA2C
We also apply the proposed advantage clipping to multi-agent advantage actor-critic (MAA2C) (Papoudakis et al., 2020), dubbed OptiMAA2C. Figure 7 demonstrates that OptiMAA2C improves MAA2C similarly to how OptiMAPPO improves MAPPO.

### v3

@@ Abstract
Relative overgeneralization (RO) occurs in cooperative multi-agent learning tasks when agents converge towards a suboptimal joint policy due to overfitting to suboptimal behaviors of other agents. No methods have been proposed for addressing RO in multi-agent policy gradient (MAPG) methods although these methods produce state-of-the-art results. To address this gap, we propose a general, yet simple, framework to enable optimistic updates in MAPG methods that alleviate the RO problem. Our approach involves clipping the advantage to eliminate negative values, thereby facilitating optimistic updates in MAPG. The optimism prevents individual agents from quickly converging to a local optimum. Additionally, we provide a formal analysis to show that the proposed method retains optimality at a fixed point. In extensive evaluations on a diverse set of tasks including the Multi-agent MuJoCo and Overcooked benchmarks, our method outperforms strong baselines on 13 out of 19 tested tasks and matches the performance on the rest.
@@ 1 Introduction
Multi-agent reinforcement learning (MARL) is a promising approach for many cooperative multi-agent decision making applications, such as those found in robotics and wireless networking (Busoniu et al., 2008). However, despite recent success on increasingly complex tasks (Vinyals et al., 2019; Yu et al., 2022), these methods can still fail on simple two-player matrix games as shown in Figure 1. The underlying problem is that in cooperative tasks, agents may converge to a suboptimal joint policy when updating their individual policies based on data generated by other agents’ policies which have not converged yet. This phenomenon is called relative overgeneralization (Wiegand, 2004) (RO) and has been widely studied in tabular matrix games (Claus & Boutilier, 1998; Lauer & Riedmiller, 2000; Panait et al., 2006) but remains an open problem in state-of-the-art MARL methods (De Witt et al., 2020; Yu et al., 2022; Kuba et al., 2022).
@@
The cause for RO can be understood intuitively. In a cooperative multi-agent system, the common reward derives from the joint actions. From the perspective of a single agent, an optimal individual action may still incur low joint reward due to non-cooperative behaviors of other agents. This is common in cooperative tasks when iteratively optimizing individual policies, especially at the beginning of training since individual agents have not learned to act properly to cooperate with others. It is particularly challenging in tasks with a large penalty for incorrect joint actions, such as the pitfalls in the climbing matrix game in Figure 1 leading often to agents that prefer a suboptimal joint policy.
@@
Existing techniques to overcome the pathology share the idea of applying optimistic updating in Q-learning with different strategies to control the degree of optimism (Lauer & Riedmiller, 2000) such as in hysteretic Q-learning (Matignon et al., 2007) or in lenient agents (Panait et al., 2006). However, these methods are designed based on Q-learning and only tested on matrix games (Matignon et al., 2012) with tabular Q representation. Although optimism successfully helps overcome RO problem and converges to global optima in tabular tasks, it could amplify the overestimation problem when combined with DQN (Van Hasselt et al., 2016) with function approximation (Omidshafiei et al., 2017; Palmer et al., 2018; Rashid et al., 2020a), as verified in our experiments. On the other hand, recent multi-agent policy gradient (MAPG) methods have achieved state-of-the-art performance on popular MARL benchmarks (De Witt et al., 2020; Yu et al., 2022; Sun et al., 2023; Wang et al., 2023) but still suffer from the RO problem, converging to a low value local optimum in simple matrix games. The above drawbacks motivate this work to investigate whether optimism can be applied to MAPG methods and whether it can boost performance further by mitigating the RO problem.
@@
Our contribution is threefold: (1) To our knowledge, we are the first to investigate the application of optimism in MAPG methods. We propose a general, yet simple, framework to incorporate optimism into the policy gradient computation in MAPG. Specifically, we propose to clip the advantage values when updating the policy. For completeness, we extend the framework to include a hyperparameter to control the degree of optimism, resulting in a Leaky ReLU function (Maas et al., 2013) to reshape the advantage values. (2) We provide a form analysis to show that the proposed method retains optimality at a fixed point. (3) In our experiments111Source Code: https://github.com/wenshuaizhao/optimappo, the proposed OptiMAPPO algorithm successfully learns global optima in matrix games and outperforms both recent state-of-the-art MAPG methods MAPPO (Yu et al., 2022), HAPPO, HATRPO (Kuba et al., 2022), and existing optimistic methods in complex domains.
@@ 2 Related work
In this section, we discuss classic optimistic methods and optimistic DQN based approaches. We also discuss recent MAPG methods that yield state-of-the-art performance on common benchmarks. Optimistic Thompson sampling (O-TS) (Hu et al., 2023) is also discussed as it utilizes a similar clipping technique to improve exploration for stochastic bandits. General multi-agent exploration methods are related and introduced. For completeness, we further discuss advantage shaping in single-agent settings.
@@
To the best of our knowledge, distributed Q-learning (Lauer & Riedmiller, 2000) proposes the first optimistic updating in independent Q-learning, where -values are only updated when the value increases. Lauer & Riedmiller (2000) also provides brief proof that the proposed optimism-based method converges to the global optimum in deterministic environments. To handle stochastic settings, hysteretic Q-learning (Matignon et al., 2007) adjusts the degree of optimism by setting different learning rates for the Q values of actions with different rewards. The frequency maximum Q value heuristic (FMQ) (Kapetanakis & Kudenko, 2002) considers changing the action selection strategies during exploration instead of modifying the updating of Q values. Lenient agent (Panait et al., 2006; Wei & Luke, 2016) employs more heuristics to finetune the degree of optimism by initially adopting an optimistic disposition and gradually transforming into average reward learners.
@@
Recently, several works have extended optimistic methods and lenient agents to Deep Q-learning (DQN). Dec-HDRQN (Omidshafiei et al., 2017) and lenient-DQN (Palmer et al., 2018) apply hysteretic Q-learning and leniency to DQN, respectively. Optimistic methods have also been combined with value decomposition methods. Weighted QMIX (Rashid et al., 2020a) uses a higher weight to update the Q-values of joint actions with high rewards and a lower weight to update the values of suboptimal actions. FACMAC (Peng et al., 2021) improves MADDPG (Lowe et al., 2017) by taking actions from other agents’ newest policies while computing the Q values, which can also be regarded as optimistic updating as it is natural to suppose the newer policies of other agents would generate better joint Q-values. Even though the optimism has been applied to DQN-based methods, in common benchmarks they are usually outperformed by recent MAPG methods (Yu et al., 2022). This unsatisfying performance of optimistic DQN methods can be attributed to the side effect of overestimation of Q-values, which we also empirically verify in our experiments.
@@
COMA (Foerster et al., 2018) is an early MAPG method in parallel with value decomposition methods (Sunehag et al., 2017; Rashid et al., 2020b). It learns a centralized on-policy function and utilizes it to compute the individual advantage for each agent to update the policy. However, until IPPO and MAPPO (De Witt et al., 2020; Yu et al., 2022) which directly apply single-agent PPO (Schulman et al., 2017) into multi-agent learning, MAPG methods show significant success on popular benchmarks. The strong performance of these methods might be credited to the property that individual trust region constraint in IPPO/MAPPO can still lead to a centralized trust region as in single-agent PPO (Sun et al., 2023). Nonetheless, HAPPO/HATRPO (Kuba et al., 2022) and A2PO (Wang et al., 2023) further improve the monotonic improvement bound by enforcing joint and individual trust region, while with the cost of sequentially updating each agent. However, even with a strong performance on popular benchmarks, these methods don’t explicitly consider the relative overgeneralization problem during multi-agent learning and can still converge to a suboptimal joint policy.
@@
Since RO can be seen as a special case of the general exploration problem we discuss below general MARL exploration methods. Similar to single-agent exploration, Rashid et al. (2020b) and Hu et al. (2021) use noise for exploration in multi-agent learning. However, as shown by (Mahajan et al., 2019) noise-based exploration can result in suboptimal policies. For coordinated exploration, multi-agent variational exploration (Mahajan et al., 2019) conditions the joint Q-value on a latent state. Jaques et al. (2019) maximizes the mutual information between agent behaviors for coordination. This may nevertheless still lead to a sub-optimal joint strategy (Li et al., 2022). Zhao et al. (2023) propose conditionally optimistic exploration (COE) which augments agents’ Q-values by an optimistic bonus based on a global state-action visitation count of preceding agents. However, COE is designed for discrete states and actions and is difficult to scale up to complex tasks with continuous state and action spaces. Cooperative multi-agent exploration (CMAE) (Liu et al., 2021) only counts the visitations of states in a restricted state space to learn an additional exploration policy. However, the restricted space selection is hard to scale up and thus CMAE fails to show performance improvement on widely used MARL benchmarks. Our method aims to mitigate the specific RO problem with state-of-the-art MAPG methods in order to further boost their performance on complex tasks.
@@ 3.1 Problem Formulation
We mainly study the fully cooperative multi-agent sequential decision-making tasks which can be formulated as a multi-agent Markov decision process (MMDP) (Boutilier, 1996) consisting of a tuple , where is the set of agents. At time step of Dec-MDP, each agent observes the full state in the state space of the environment, performs an action in the action space from its policy . The joint policy consists of all the individual policies . The environment takes the joint action of all agents , changes its state following the dynamics function and generates a common reward for all the agents. is a reward discount factor. The agents learn their individual policies and maximize the expected return: . Since we train independent policy networks for each agent, we use and interchangeably as they all include the full state information in our experiments.
@@ 4.1 Optimistic MAPPO
The advantage is usually estimated by the generalized advantage estimator (GAE) (Schulman et al., 2016) defined as
@@
In order to overcome the relative overgeneralization problem in MAPG methods, the proposed optimistic MAPPO (OptiMAPPO) applies a clipping operation to reshape the estimated advantages (Schulman et al., 2016). OptiMAPPO optimizes the agents’ policies by maximizing the following new objective
@@
The proposed clipping operation can be seen as a special case of a Leaky ReLU (LR) operation of advantage values where there is a hyperparameter to control the degree of optimism, . Our clipping operation is the case when , while recovers the original MAPG methods. In our experiments, we find that the performance is improved more while setting lower , i.e. higher optimism. However, we argue that such an extension could be beneficial in stochastic reward environments as discussed in hysteretic Q-learning (Matignon et al., 2007) and lenient agents methods (Palmer et al., 2018; Matignon et al., 2012). The extension allows us to control the degree of optimism in a finer granularity. We leave the study in stochastic environments as future work and this paper is primarily focused on the first to investigate the effectiveness of optimism in MAPG methods.
@@ 5.1 Repeated Matrix Games
Even though with small state and action spaces, the climbing and penalty matrix games (Claus & Boutilier, 1998), as shown in Figure 1, are usually hard to obtain the optimal joint solution without explicitly overcoming the relative overgeneralization problem. The matrix games have two agents which select the column and row index of the matrix respectively. The goal is to select the correct row and column index to obtain the maximal element of a matrix. In the penalty game, is the penalty term and we evaluate for . The lower value of , the harder for agents to identify the optimal policy due to the growing risk of penalty. Following (Papoudakis et al., 2021), we set constant observation and the episode length of repeated games as .
@@ 5.6 Does Q-learning Based Optimism Work?
We empirically analyze whether Q-learning based optimism can boost performance in the Overcooked tasks using hysteretic Q-learning (Omidshafiei et al., 2017; Palmer et al., 2018). Figure 5 shows average Q values during training with different degrees of optimism. When decreases excessively, i.e. optimism increases too much, performance decreases. In both tasks shown in Figure 5, the highest optimism estimates the highest Q-values while showing the worst performance.
@@ 6 Limitation and Conclusion
Our optimistic updating approach yields state-of-the-art performance. However, as with other optimistic updating methods (Lauer & Riedmiller, 2000; Matignon et al., 2007), optimism can lead to a sub-optima when misleading stochastic rewards exist. The proposed Leaky ReLU extension allows further adaptive adjustment of the optimism degree to balance between optimism and neutrality and may allow to reduce the severity of stochastic rewards but this requires future investigation. In addition, lenient agents (Panait et al., 2006; Wei & Luke, 2016; Palmer et al., 2018) provide a set of heuristic techniques to adapt the degree of optimism, applicable to our method, to mitigate the problem with stochastic rewards. We leave this also as future work.
@@ C.2 Key Hyper-Parameters
In all the tasks of MA-MuJoCo, we use the same hyperparameters listed in Table 5. The implementation is based on the HAPPO (Kuba et al., 2022) codebase, and the other hyperparameters are the default.
@@ Appendix D More Results of Popular Deep MARL methods on Matrix Games
In order to show the RO problem in existing MARL methods, we cite the following results of common deep MARL methods on the penalty and climbing games from the benchmarking paper (Papoudakis et al., 2021). As shown in Table 7, popular MARL methods without explicitly considering the RO problem fail to solve the matrix games.
@@ Appendix E Results of Optimistic MAA2C
We also apply the proposed advantage clipping to multi-agent advantage actor-critic (MAA2C) (Papoudakis et al., 2021), dubbed OptiMAA2C. Figure 7 demonstrates that OptiMAA2C improves MAA2C similarly to how OptiMAPPO improves MAPPO.